Waqas wishes to distribute 128 bananas and also 176 apples equally among a certain number of children.find the highest number of children who can get the fruit in this way
1
Expert's answer
2012-08-23T11:33:57-0400
Equally, we need to find the greatest common divisor of 128 and 176:

gcd(128,176)
= 2*gcd(64,88)
= 4*gcd(32,44)
= 8*gcd(16,22)
= 16*gcd(8,11)
= |11 is a prime number|
= 16*1 = 16.

So, the highest number of children who can get the fruit is 16.